Find the standard of the line.
(-5,8) (4,-3)

Answer by Adrish(44)

SOLUTIONS:
Standard Equation Is-
Ax+By=C
Given:(-5,8) and (4,-3)
From here we can find the slope
Slope(m)= Rise/Run
=(-3-8)/(4+5)
=-11/9
Equation will be
y-y1=m(x-x1)
y-8=-11/9(x+5)
y-8=-11/9x - 55/9
11/9x + y = -55/9 +8
11/9x + y = 17/9